# DiffLM: Controllable Synthetic Data Generation via Diffusion Language Models

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their knowledge and generative capabilities, leading to a surge of interest in leveraging LLMs for high-quality data synthesis. However, synthetic data generation via prompting LLMs remains challenging due to LLMs’ limited understanding of target data distributions and the complexity of prompt engineering, especially for structured formatted data. To address these issues, we introduce DiffLM, a controllable data synthesis framework based on variational autoencoder (VAE), which further (1) leverages diffusion models to reserve more information of original distribution and format structure in the learned latent distribution and (2) decouples the learning of target distribution knowledge from the LLM’s generative objectives via a plug-and-play latent feature injection module. As we observed significant discrepancies between the VAE’s latent representations and the real data distribution, the latent diffusion module is introduced into our framework to learn a fully expressive latent distribution. Evaluations on seven real-world datasets with structured formatted data (i.e., Tabular, Code and Tool data) demonstrate that DiffLM generates high-quality data, with performance on downstream tasks surpassing that of real data by 2%–7% in certain cases. The data and code will be publicly available upon completion of internal review.
@@ 1 Introduction
Despite the progress achieved, generating high-quality synthetic textual data using LLMs remains challenging, particularly for structured data (Josifoski et al., 2023; Li et al., 2022). First, LLMs often lack a global understanding of the target data distribution when generating synthetic data. Even after fine-tuning, it is difficult to inject information about complex and varied distributions into current LLM architectures, often resulting in outputs with low diversity and instances of data copying (Wu et al., 2024; Yu et al., 2023). Moreover, existing LLM-based synthetic data generation methods typically involve complex pipelines and post-processing mechanisms, such as prompt engineering, multi-agent frameworks, and iterative sampling (Dekoninck et al., 2024; Wu et al., 2024). These complexities hinder the rapid adaptation of LLMs to new tasks, limiting their utility in dynamic research and industrial scenario. Concurrently, the remarkable performance of variational autoencoders (VAEs) and diffusion models in image synthesis tasks (Betker et al., 2023; Rombach et al., 2022) has spurred interest in adapting these techniques to other modalities (Borisov et al., 2023; Li et al., 2022; Gong et al., 2023). Although some works have introduced latent spaces into language models for simple tasks like style transfer or topic generation (Yang & Klein, 2021; Li et al., 2022), our preliminary experiments indicate that directly applying the latent distributions learned by VAEs often results in outputs that are unrelated to the real data. This challenges the direct application of these methods in more complex scenarios for synthetic data generation.
@@
To address these challenges, we propose DiffLM, a novel framework that leverages a plug-and-play latent space to provide data distribution information for LLMs during data generation. First, to decouple the learning of real data distributions from the LLM’s training objectives, we develop a latent space using a VAE to capture external information, mapping samples from the real dataset to latent vectors. However, we observed that simply sampling from a Gaussian distribution obtained from naive VAE that cannot generate realistic results. To overcome the poor quality of data generated by sampling from VAE, we employ a latent diffusion method that linearly adds noise to the latent space over time. A denoising network is then trained to learn these noises in the reverse process, reducing efficiency loss in data synthesis due to sampling failures. Finally, we design a soft prompting method to inject latent features into the LLM decoding process, resulting in controllable, high-quality synthetic data. We evaluate our method on seven real-world structured formatted datasets, ranging from simple table synthesis to more complex code and tool synthesis tasks. Experiments demonstrate that DiffLM can generate realistic results, and ablation studies confirm the effectiveness of each component in our proposed method. The contributions of this paper are threefold:
@@
  • High-Quality Synthetic Data Generation: Based on our observations, the meticulously designed VAE and diffusion structures effectively model the distribution of real data, enabling the generation of high-quality synthetic data. In all tasks, the quality of the generated data is comparable to or even surpasses that of the real data.
@@ 2 Related Works
Large Language Models in Data Synthesis.
@@
Latent Variable Models in Text Generation.
Latent variable models have made significant advances in computer vision in recent years (Yu et al., 2022; Gu et al., 2022; Luo et al., 2023; Gulrajani et al., 2017), achieving high-quality generation results, flexibility and effectiveness, as well as robustness to noise perturbations. In particular, latent diffusion models, such as DALL-E (Betker et al., 2023) and Stable Diffusion (Rombach et al., 2022), operate their diffusion processes in a latent space rather than directly in data space, enabling a near-optimal balance between generation quality and computational efficiency. In text generation, several works (Wiseman et al., 2018; Ding & Gimpel, 2019; Li et al., 2022; Gu et al., 2023; Borisov et al., 2023) have attempted to combine latent spaces with language models to accomplish tasks such as sentence representation, text style transfer, and dataset augmentation. Additionally, some studies have explored the use of diffusion for plug-and-play controllable generation (Li et al., 2022; Gong et al., 2023), aiming to steer the outputs of pre-trained language model using auxiliary modules. While these works share a similar perspective with ours, we tackle a more challenging scenario of structured data synthesis and thoroughly investigate multiple methods of latent knowledge injection. To the best of our knowledge, our work is the first to combine VAEs and diffusion models with LLMs for high-quality data synthesis.
@@ 3 Methodology
Figure 1 illustrates the main pipeline of our proposed DiffLM. First, we define an encoder to map discrete data into a continuous latent space (Section 3.2). Second, although text features are extracted and compressed, vanilla latent embeddings in VAEs often lead to decoding failures due to underutilized or empty regions in the latent space. To address this issue, we adopt a latent diffusion model (Section 3.3). Finally, to incorporate the prior knowledge into the decoding stage of LLMs, we propose a novel soft prompt injection method to steer the decoding process (Section 3.4).
@@ 3.2 VAE-based Representation Learning
Feature Encoding: In standard VAEs, an encoder is typically employed to map input data into a latent space. Given structured text data $s_{i}$, we utilize a learnable Transformer-based pre-trained language model (Vaswani et al., 2017; Devlin et al., 2019; Raffel et al., 2020) to obtain the representation vector $x_{i}\in\mathbb{R}^{d\times 2}$, which can be split into the mean and variance. Using the re-parameterization trick (Kingma & Welling, 2014), we then obtain the latent feature $z\in\mathbb{R}^{d}$:
@@
LLM Decoding: After generating the latent feature $z$, we employ a frozen-parameter LLM to reconstruct the input text $s$ in a causal language modeling manner. The rationale for freezing the LLM parameters is to avoid retraining and to preserve its general knowledge and capabilities. Consequently, aligning the two different modalities, whereas the latent space and the LLM input space, presents a significant challenge. To address this, we propose a novel latent feature injector using soft prompting and design a corresponding injector network; specific details are provided in Section 3.4.
VAE Training Objective: The VAE model is typically trained using the Evidence Lower Bound (ELBO) loss function. Following previous work (Burgess et al., 2018), we adopt the $\beta$-VAE training strategy, which introduces a weighting parameter $\beta$ to control the contribution of the KL divergence loss in the total loss function. Specifically, when $\beta=0$, the model reduces to a standard autoencoder. For $\beta>0$, the KL constraint encourages learning a smoother latent space:
@@ 3.3 Latent Space Denoising
$dz_{t}$ $=-\dot{\sigma}(t)\sigma(t)\nabla_{z_{t}}\log{p(z_{t})}dt+\sqrt{2\dot{\sigma}(t)\sigma(t)}d\omega_{t},$ (6)
@@
In forward process Eq.5, $z_{t}$ is the latent variable at time $t$, and $\sigma(t)$ is a time-dependent noise scale function. As for backward process Eq.6, $\dot{\sigma}(t)$ stands for the time derivative of $\sigma(t)$, and $\nabla_{z_{t}}\log{p(z_{t})}$ is the gradient of the log probability density with respect to $z_{t}$, also known as the score function, and $d\omega_{t}$ is an increment of the Wiener process (standard Brownian motion). For diffusion model training loss Eq.7, $\epsilon_{\theta}(z_{t},t)$ is the neural network that predicts the noise $\epsilon$ given $z_{t}$ and $t$. The detailed description for diffusion model could be found in Appendix A.1.
3.4 Latent Feature Injection
After constructing a latent space that captures the true data distribution, two challenges remain: 1) Aligning latent space with LLM’s input space. How can the decoding LLM process the latent vector $z$ to steer a powerful language model for realistic data generation? 2) Seamless Integration with LLM Knowledge: How can we integrate external information without disrupting the LLM’s internal knowledge? Motivated by adapter training methods in LLM fine-tuning (Lester et al., 2021; Li & Liang, 2021; Houlsby et al., 2019; Liu et al., 2023a), we consider the soft prompt latent injection approach to incorporate $z$ into LLM decoding without training the model weights. Specifically, after obtaining the latent representation $z$, we use an upper MLP to map it into $k$ soft prompt token embeddings, denoted as $\mathbf{H_{\text{latent}}}\in\mathbb{R}^{k\times d}$. These soft embeddings serve as a steering vector, which is concatenated before the $<$BOS$>$ token to assist the LLM in decoding. The detailed process is illustrated in Figure 4. We also conduct ablation experiments in Section 5.1 with the other two injection methods proposed by Li et al. (2020), which validated that our methods obtain the best reconstruction loss and downstream task performance.
4 Experiments
In this section, we evaluate the generation quality of the DiffLM method on multiple public benchmarks across three tasks: 1) Tabular Data Generation: We compare DiffLM with SoTA tabular generation algorithms, demonstrating its strong capability in structured data generation. 2) Code Generation: DiffLM showcases the ability to integrate structured data priors with its internal knowledge. The results on synthetic data are even better than real ones. 3) Tool Generation: DiffLM can quickly adapt to complex function call scenarios, highlighting its flexibility and adaptability.
Table 1: Performance of downstream tasks using generated tabular data. We evaluate the quality from: performance in machine learning efficiency (MLE) task, and column-wise distribution density estimation ($\bm{\rho}$) task. $\uparrow$,$\downarrow$ , ↓ indicate that higher (or lower) metrics correspond to better performance. Boldface indicates DiffLM surpasses the SoTA model based on language models. Red Boldface denotes DiffLM exceeds the MLE performance achieved using real data.
@@ 4 Experiments
DiffLM 0.894 9.16 0.793 9.33 0.910 7.04 0.912 14.43 0.717 6.05
4.1 Tabular Data Generation
Baselines. We selected a comprehensive set of classic and SoTA tabular data generation models with diverse architectures for comparison. First, we consider the performance on real data as the upper bound for evaluation. Secondly, we included the classic method, synthetic minority over-sampling technique (SMOTE) (Chawla et al., 2002), which generates new synthetic data patterns by performing linear interpolation between minority class samples and their $k$ nearest neighbors. Additionally, for neural network-based tabular generation algorithms, we considered six baseline models across different architectures: 1) GAN-based models: CTGAN (Xu et al., 2019); 2) VAE-based models: TVAE (Xu et al., 2019), GOGGLE (Liu et al., 2023b); 3) Diffusion-based models: Codi (Lee et al., 2023), TabSyn (Zhang et al., 2024); 4) LLM-based: GReaT (Borisov et al., 2023), which attempts to fine-tune a GPT-2 (Radford et al., 2019) for table synthesis. It is worth noting that we compare with the current strongest generative models not to merely outperform them in tabular generation but to demonstrate that our flexible DiffLM architecture can achieve comparable performance while offering additional advantages.
Evaluation. Table 1 presents the quality assessment results of the generated data. For different tabular datasets, we train a XGBoost classifier or a regressor using the synthetic data to predict the label column values, using AUC and RMSE to evaluate the accuracy, respectively. From the results, DiffLM outperforms the current language-model-based SoTA (GReaT model) on most datasets. Notably, on the Default dataset, the prediction accuracy using DiffLM’s synthetic data surpasses that obtained by training on real data. This suggests that DiffLM’s approach of integrating the real data distribution with its own learned knowledge can provide richer information for downstream tasks while preserving the original data structure. In other words, the synthetic data generated by DiffLM contains additional knowledge compared to real data, which is challenging to achieve with previous methods. Moreover, our generated results achieve performance comparable to prior methods in column-wise distribution density estimation. Although the TabSyn method attains superior performance on several datasets, it should be noted that our approach focuses on general, pluggable generation control for large language model, rather than training data synthesis models from scratch for specific domains. Despite this, in tabular data generation, our method’s performance is on par with these domain-specific methods.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their knowledge and generative capabilities, leading to a surge of interest in leveraging LLMs for high-quality data synthesis. However, synthetic data generation via prompting LLMs remains challenging due to LLMs’ limited understanding of target data distributions and the complexity of prompt engineering, especially for structured formatted data. To address these issues, we introduce DiffLM, a controllable data synthesis framework based on variational autoencoder (VAE), which further (1) leverages diffusion models to reserve more information of original distribution and format structure in the learned latent distribution and (2) decouples the learning of target distribution knowledge from the LLM’s generative objectives via a plug-and-play latent feature injection module. As we observed significant discrepancies between the VAE’s latent representations and the real data distribution, the latent diffusion module is introduced into our framework to learn a fully expressive latent distribution. Evaluations on seven real-world datasets with structured formatted data (i.e., Tabular, Code, and Tool data) demonstrate that DiffLM generates high-quality data, with performance on downstream tasks surpassing that of real data by 2%–7% in certain cases. Data and code are available at https://github.com/bytedance/DiffLM.
DiffLM: Controllable Synthetic Data Generation via Diffusion Language Models
Ying Zhou1,2thanks: Equal contribution., Xinyao Wang3∗thanks: Work done before joining Amazon AGI., Yulei Niu3‡, Yaojie Shen1,2, Lexin Tang3, Fan Chen3, Ben He1,2‡, Le Sun2, Longyin Wen3thanks: Corresponding authors. 1University of Chinese Academy of Sciences, Beijing, China 2Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences, Beijing, China 3ByteDance Inc., San Jose, USA zhouying20@mails.ucas.ac.cn
@@ 1 Introduction
Despite the progress achieved, generating high-quality synthetic textual data using LLMs remains challenging, particularly for structured data (Josifoski et al., 2023; Li et al., 2022). First, LLMs often lack a global understanding of the target data distribution when generating synthetic data. Even after fine-tuning, it is difficult to inject information about complex and varied distributions into current LLM architectures, often resulting in outputs with low diversity and instances of data copying (Wu et al., 2024; Yu et al., 2023). Moreover, existing LLM-based synthetic data generation methods typically involve complex pipelines and post-processing mechanisms, such as prompt engineering, multi-agent frameworks, and iterative sampling (Dekoninck et al., 2024; Wu et al., 2024). These complexities hinder the rapid adaptation of LLMs to new tasks, limiting their utility in dynamic research and industrial scenario. Concurrently, the remarkable performance of variational autoencoders (VAEs) and diffusion models in image synthesis tasks (Betker et al., 2023; Rombach et al., 2022) has spurred interest in adapting these techniques to other modalities (Borisov et al., 2023; Li et al., 2022; Gong et al., 2023). Although some works have introduced latent spaces into language models for simple tasks like style transfer or topic generation (Yang and Klein, 2021; Li et al., 2022), our preliminary experiments indicate that directly applying the latent distributions learned by VAEs often results in outputs that are unrelated to the real data. Similar issues also have been addressed in prior works (Amani et al., 2024; Havrylov and Titov, 2020; Bowman et al., 2016). This challenges the direct application of these methods in more complex scenarios for synthetic data generation.
@@
To address these challenges, we propose DiffLM, a novel framework that leverages a plug-and-play latent space to provide data distribution information for LLMs during data generation. First, to decouple the learning of real data distributions from the LLM’s training objectives, we develop a latent space using a VAE model to capture external information, mapping samples from the real dataset to latent vectors. However, we observed that sampling points from a Gaussian distribution obtained from naive VAE that cannot generate realistic results. To overcome the poor quality of data generated by sampling from VAE, we employ a latent diffusion method that linearly adds noise to the latent space over time. A denoising network is then trained to learn these noises in the reverse process, reducing efficiency loss in data synthesis due to sampling failures. Finally, we design a soft prompting method to inject latent features into the LLM decoding process, resulting in controllable, high-quality synthetic data. We evaluate our method on seven real-world structured formatted datasets, ranging from relatively simple table synthesis to more complex code and tool synthesis tasks. Experiments demonstrate that DiffLM can generate high-quality results, and ablation studies confirm the effectiveness of each component in our proposed method. The contributions of this paper are threefold:
@@
  • High-Quality Synthetic Data: Based on our observations, the meticulously designed VAE and diffusion structures effectively model the distribution of real data, enabling the generation of high-quality synthetic data. In all tasks, the quality of the generated data is comparable to or even surpasses that of the real data.
@@ 2 Related Works
Large Language Models for Data Synthesis.
@@
Latent Variable Models for Text Generation.
Latent variable models have made significant advances in computer vision in recent years (Yu et al., 2022a; Gu et al., 2022; Luo et al., 2023a; Gulrajani et al., 2017), achieving high-quality generation results, flexibility and effectiveness, as well as robustness to noise perturbations. In particular, latent diffusion models, such as DALL-E (Betker et al., 2023) and Stable Diffusion (Rombach et al., 2022), operate their diffusion processes in a latent space rather than directly in data space, enabling a near-optimal balance between generation quality and computational efficiency. In text generation, several works (Bowman et al., 2016; Kaiser and Bengio, 2018; Havrylov and Titov, 2020; Li et al., 2022; Gu et al., 2023; Borisov et al., 2023; Amani et al., 2024) have attempted to combine latent spaces with language models to accomplish tasks such as language modeling (Lin et al., 2023; Reid et al., 2023; Zhang et al., 2023; Yuan et al., 2024), unconditional text generation (Yu et al., 2022b; Chen et al., 2023) and control text generation (Austin et al., 2021a; He et al., 2023). Additionally, some studies have explored the use of diffusion for plug-and-play controllable generation (Li et al., 2022; Gong et al., 2023), aiming to steer the outputs of pre-trained language model using auxiliary modules. While these works share a similar perspective with ours, we tackle a more challenging scenario of structured data synthesis. To the best of our knowledge, our work is the first to combine VAEs and denoising diffusion models with large language models for high-quality data synthesis.
@@ 3 Methodology
Figure 1 illustrates the main pipeline of our proposed DiffLM. First, we define an encoder to map discrete text into a continuous latent space (Section 3.2). Second, although the features of the data are extracted and compressed, conventional latent embeddings in text VAEs often lead to decoding failures due to underutilized or empty regions in the latent space. To address this issue, we train a diffusion model on the latent space (Section 3.3). Finally, to incorporate encoded prior knowledge into the decoding stage of large language models, we propose a novel soft prompt injection method to steer the decoding process (Section 3.4).
@@ 3.2 VAE-based Representation Learning
Feature Encoding: In standard VAEs, an encoder is typically employed to map input data into a latent space. Given structured text data $s_{i}$, we utilize a learnable Transformer-based pre-trained language model (Vaswani et al., 2017; Devlin et al., 2019; Raffel et al., 2020) to obtain the representation vector $x_{i}\in\mathbb{R}^{d\times 2}$, which can be split into the mean and variance. Using the re-parameterization trick (Kingma and Welling, 2014), we then obtain the latent feature $z\in\mathbb{R}^{d}$:
@@
LLM Decoding: After generating the latent feature $z$, we employ a frozen-parameter LLM to reconstruct the input text $s$ in a causal language modeling manner. The rationale for freezing the LLM parameters is to avoid retraining and to preserve its general knowledge and reasoning capabilities. Consequently, aligning the two different modalities, whereas the latent space and the LLM input space, presents a significant challenge. To address this, we propose a novel latent feature injector using soft prompting and design a corresponding injector network; specific details are provided in Section 3.4.
VAE Model Training Objective: The VAE model is typically trained using the Evidence Lower Bound (ELBO) loss function. Following previous work (Burgess et al., 2018), we adopt the $\beta$-VAE training strategy (Higgins et al., 2017), which introduces a weighting parameter $\beta$ to control the contribution of the KL divergence loss in the total loss function. Specifically, when $\beta=0$, the model reduces to a standard autoencoder. For $\beta>0$, the KL constraint encourages learning a smoother latent space:
@@ 3.3 Latent Space Denoising
$dz_{t}=$ $-\dot{\sigma}(t)\sigma(t)\nabla_{z_{t}}\log{p(z_{t})}dt$ (6)
$+\sqrt{2\dot{\sigma}(t)\sigma(t)}d\omega_{t}.$
In forward process Eq.5, $z_{t}$ is the latent variable at time $t$, and $\sigma(t)$ is a time-dependent noise scale function. As for backward process Eq.6, $\dot{\sigma}(t)$ stands for the time derivative of $\sigma(t)$, and $\nabla_{z_{t}}\log{p(z_{t})}$ is the gradient of the log probability density with respect to $z_{t}$, also known as the score function, and $d\omega_{t}$ is an increment of the Wiener process (standard Brownian motion). As for diffusion model training loss Eq.7, $\epsilon_{\theta}(z_{t},t)$ is the neural network that predicts the noise $\epsilon$ given $z_{t}$ and $t$.
@@
The detailed description for diffusion model could be found in Appendix A.2.
Table 1: Performance of downstream tasks using generated tabular data. We evaluate the quality from: performance in machine learning efficiency (MLE) task, and column-wise distribution density estimation ($\bm{\rho}$) task. $\uparrow$,$\downarrow$ , ↓ indicate that higher (or lower) metrics correspond to better performance. Boldface indicates DiffLM surpasses the SoTA model based on language models. Red Boldface denotes DiffLM exceeds the performance achieved using real data.
@@
DiffLM (ours) 0.906 9.74 0.794 9.06 0.917 7.53 0.915 10.07 0.696 6.35
3.4 Latent Feature Injection
After constructing a latent space that captures the true data distribution, two challenges remain: 1) Aligning latent space with LLM’s input space. How can the decoding LLM process the latent vector $z$ to steer a powerful language model for realistic data generation? 2) Seamless Integration with LLM Knowledge: How can we integrate external information without disrupting the LLM’s internal knowledge? Motivated by adapter training methods in LLM fine-tuning (Lester et al., 2021; Li and Liang, 2021; Houlsby et al., 2019; Liu et al., 2023a), we consider the soft prompt latent injection approach to incorporate $z$ into LLM decoding without training the model weights. Specifically, after obtaining the latent representation $z$, we use an upper MLP to map it into $k$ soft prompt token embeddings, denoted as $\mathbf{H_{\text{latent}}}\in\mathbb{R}^{k\times d}$. These soft embeddings serve as a steering vector, which is concatenated before the $<$BOS$>$ token to assist the LLM in decoding. The detailed process is illustrated in Figure 4. We also conduct ablation experiments in Section 5 with the other two injection methods proposed by Li et al. (2020), which validated that our methods obtain the best reconstruction loss and downstream task performance.
4 Experiments
In this section, we evaluate the generation quality of the DiffLM method on multiple public benchmarks across three tasks: 1) Tabular Data Generation: We compare DiffLM with SoTA tabular generation algorithms, demonstrating its strong capability in structured data generation. 2) Code Generation: DiffLM showcases the ability to integrate structured data priors with its internal knowledge. The results on synthetic data are even better than real ones. 3) Tool Generation: DiffLM can quickly adapt to complex function call scenarios, highlighting its flexibility and adaptability. All experiments adopt Mistral-v0.3 as the frozen VAE decoder. Additional implementation details for reproduction can be found in Appendix B.3.
@@ 4 Experiments
4.1 Tabular Data Generation
Baselines. We selected a comprehensive set of classic and SoTA tabular data generation models with diverse architectures for comparison. First, we consider the performance on real data as the upper bound for evaluation. Secondly, we included the classic method, synthetic minority over-sampling technique (SMOTE) (Chawla et al., 2002), which generates new synthetic data patterns by performing linear interpolation between minority class samples and their $k$ nearest neighbors. Additionally, for neural network-based tabular generation algorithms, we considered six baselines across different architectures: 1) GAN-based models: CTGAN (Xu et al., 2019); 2) VAE-based models: TVAE (Xu et al., 2019), GOGGLE (Liu et al., 2023b); 3) Diffusion-based models: CoDi (Lee et al., 2023), TabSyn (Zhang et al., 2024); 4) LLM-based: 5-shot ICL-based generation, where structural information for each dataset (specifically, column names and data types for each row) is also provided. We leverage GPT-4, Mistral-v0.3-7B-instruct, and Qwen2.5-14B-instruct models for synthetic data generation. Additionally, we include GReaT (Borisov et al., 2023), which attempts to fine-tune a GPT-2 (Radford et al., 2019) for table synthesis. We also report the GReaT result with Mistral model (Jiang et al., 2023) in Appendix C.6. It is worth noting that we compare with the SoTA generative models not to merely outperform them in tabular generation but to demonstrate that our flexible DiffLM can achieve comparable performance while offering additional advantages.
Evaluation. Table 1 presents the quality assessment results of the generated data. For different tabular datasets, we train an XGBoost classifier or regressor on the synthetic data to predict label values, evaluating performance using AUC and RMSE. The results indicate that the performance of the LLM-based synthetic data generation methods leveraging ICL is generally lower; even GPT-4, the best-performing ICL-based model, consistently achieves inferior results across all datasets compared to the fine-tuned GReaT model. However, our proposed DiffLM method consistently outperforms the current LLM-based SoTA GReaT on most datasets. Notably, on the Default dataset, the prediction accuracy using DiffLM’s synthetic data surpasses that obtained by training on real data. This suggests that DiffLM’s approach of integrating the real data distribution with its own learned knowledge can provide richer information for downstream tasks while preserving the original data structure. In other words, the synthetic data generated by DiffLM contains additional knowledge compared to real data, which is challenging to achieve with previous methods. Moreover, our generated results achieve performance comparable to prior methods in column-wise distribution density estimation. Although the TabSyn method attains superior performance on several datasets, it should be noted that our approach focuses on general, pluggable generation control for large language model, rather than training data synthesis models from scratch for specific domains. Despite this, in tabular data generation, our method’s performance is on par with these domain-specific methods.